40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Feb 14 02:09:55 2014, ericq, Update, Electronics, Transmon QPD whitening qpdWhitening.pdf
    Reply  Mon Feb 17 20:35:19 2014, ericq, Update, Electronics, Transmon QPD whitening qpdY2whit.pdfETMY_QPD_whitening.zip
       Reply  Tue Jul 8 22:53:02 2014, Jenne, Rana, Update, Electronics, Transmon QPD / whitening 
Message ID: 10157     Entry time: Tue Jul 8 22:53:02 2014     In reply to: 9642
Author: Jenne, Rana 
Type: Update 
Category: Electronics 
Subject: Transmon QPD / whitening 

We need to work farther on checking out the end transmission QPD electronics situation. 

In bullet-point form, we need to:

* Ensure that the Weiss QPD head modifications have been made on these diodes.  (cf. Rai W's LLO elogs on QPDs)

* Ensure that the QPD biases are somewhere in the range of 10-15V, not the old 100V.  (Because we only need HV to make the capacitance low for RF use. Low voltage means less power dissipation in the head)

* Ensure the Rana/Rob modifications have been propagated to the whitening boards, so that we have full dynamic range.  (Steve is looking for the marked up paper schematics)

* Replace signal path resistors with low noise metal film resistors.

* Check QPDs / whitening boards for oscillation (with a scope probe), ensure that we chose an appropriate analog gain.

 

In thinking about the transimpedances that we want, we thought about the current that we expect.  We should get about 100 mW of light transmitted through the ETMs when we have full IFO lock.  There is a 50/50 BS to split the light between the QPD and the Thorlabs transmission diode, so we have about 50 mW incident on the QPDs, which is about 13 mW per quadrant.  With a sensitivity of about 0.15 Amps/Watt for silicon, this means that we're expecting to see about 2 mA of current per quadrant once we have the IFO fully resonant. We want this to correspond to about 5V, which means we want a transimpedance gain of around 2.5 kOhm. 

 

For the things that need checking, each quadrant has:

Photodiode  ------  Gain Switch 1 ----- Gain Switch 2  ------ Gain Switch 3 ------ Variable Gain Amplifier ------- Whitening stage 1 (z @ 4 Hz, p @ 40 Hz)  ------- Whitening stage 2 (z @ 4 Hz, p @ 40 Hz)

We want to check on the status of each of these switches, and whether they actually do what they say on the QPD Head screens.  Q has checked out and fixed the bit outputs for the whitening stages, but the rest still needs to be checked out.  Also note that the Switch 1, Switch 2 and Switch 3 are common to all 4 quadrants (i.e. enabling switch 1 on one quadrant enables it on all quadrants), but the variable gains and the whitening stages are individual for each quadrant.

ELOG V3.1.3-